# NUBO: A Transparent Python Package for Bayesian Optimization

### v2

@@ Abstract
NUBO, short for Newcastle University Bayesian Optimization, is a Bayesian optimization framework for optimizing expensive-to-evaluate black-box functions, such as physical experiments and computer simulators. Bayesian optimization is a cost-efficient optimization strategy that uses surrogate modeling via Gaussian processes to represent an objective function and acquisition functions to guide the selection of candidate points to approximate the global optimum of the objective function. NUBO focuses on transparency and user experience to make Bayesian optimization accessible to researchers from all disciplines. Clean and understandable code, precise references, and thorough documentation ensure transparency, while a modular and flexible design, easy-to-write syntax, and careful selection of Bayesian optimization algorithms ensure a good user experience. NUBO allows users to tailor Bayesian optimization to their problem by writing a custom optimization loop using the provided building blocks. It supports sequential single-point, parallel multi-point, and asynchronous optimization of bounded, constrained, and mixed (discrete and continuous) parameter input spaces. Only algorithms and methods extensively tested and validated to perform well are included in NUBO. This ensures that the package remains compact and does not overwhelm the user with an unnecessarily large number of options. The package is written in Python but does not require expert knowledge of Python to optimize simulators and experiments. NUBO is distributed as open-source software under the BSD 3-Clause license.
Keywords Bayesian optimization  $\cdot$ black-box optimization  $\cdot$ surrogate model  $\cdot$ Gaussian process  $\cdot$ Monte Carlo  $\cdot$ design of experiments  $\cdot$ Python
@@ 1 Introduction
The optimization of expensive black-box functions is a common problem encountered by researchers in a wide range of disciplines, such as engineering, computing, and natural sciences. These functions are characterized by an unknown or not analytically solvable mathematical expression and high evaluation costs. The principal way to gather information about a black-box function is to provide it with some inputs and observe its corresponding output. However, this process produces high costs, for example, material costs for physical experiments, computing costs for simulators, or time costs in general [1]. Many optimization algorithms, such as Adam [2], L-BFGS-B [3], and differential evolution [4], rely either on derivative information about the objective function or large numbers of function evaluations. Neither is typically feasible when working with an expensive black-box function, requiring us to search elsewhere for a cost-effective and sample-efficient alternative.
@@
Bayesian optimization takes a surrogate model-based approach to optimize expensive black-box functions in a minimum number of function evaluations. Although the genesis of Bayesian optimization can be traced back to the middle of the 20th century [5, 6, 7, 8], it gained considerable popularity in the last two decades [9, 10, 11, 1]. Recently, it has been applied to simulators and experiments in various research areas. For example, Bayesian optimization was used in the field of computational fluid dynamics to maximize the drag reduction via the active control of blowing actuators [12, 13, 14], in chemical engineering for molecular design, drug discovery, molecular modeling, electrolyte design, and additive manufacturing [15], and in computer science to fine-tune hyper-parameters of machine learning models [16] and for architecture search of neural networks [17].
@@
With NUBO, we provide an open-source implementation of Bayesian optimization aimed at researchers with expertise in disciplines other than statistics and computer science. To ensure that our target audience can understand and use Bayesian optimization to its full potential, NUBO focuses particularly on (a) transparency through clean and understandable code, precise references, and thorough documentation and (b) user experience through a modular and flexible design, easy syntax, and a careful selection of implemented algorithms. Various Python packages for Bayesian optimization exist as listed in Table 1. Most of them only support sequential single-point optimization, i.e., every point suggested by the algorithm has to be evaluated by the objective function before moving on to the next iteration. However, parallelism can be exploited in many cases to speed up the optimization process. For example, consider a simulator that can be run in parallel. Evaluating all points in parallel would save time as it would only take as long as evaluating a single point sequentially. pyGPGO [18], bayes_opt [^1] The package is also known under the name bayesian-optimization. [19], Spearmint [20], and SMAC3 [21] do not allow parallel multi-point optimization. Furthermore, Spearmint is not modular, resulting in less flexible implementations and giving the user less control when tailoring Bayesian optimization to unique research problems. To our knowledge, the closest available package to NUBO is BoTorch [22] as it also supports parallel and asynchronous optimization through the use of Monte Carlo approximations of the acquisition functions. However, compared to the lightweight implementation of NUBO, BoTorch uses a very large code base that makes code comprehension difficult, as it often requires retracing various functions and objects through many files. This can be quantified by the huge codebase represented in Table 1 as the total number of lines of code [^2] The total number of lines of code does not include comments, blank lines, and files that are irrelevant to the actual algorithms, such as examples, tests, and test functions.: NUBO implements Bayesian optimization in only 1,322 lines of code over 20 files, while BoTorch uses 38,419 lines of code—roughly 29 times more than NUBO—and spreads them between 160 files. It also provides a large number of functions and methods that enforce decisions non-expert users do not have the knowledge and experience to make. NUBO lightens this burden of the user by limiting itself to the most important methods. Table 1 also includes GPyOpt [23]; however, it is no longer maintained and has recently been archived.
@@
The number of code lines regards the underlying code bases of the packages and not the lines of code a user must write to apply Bayesian optimization. When talking about a transparent implementation, the former is a better proxy as it reflects the complexity of the whole package, that is, all functions and algorithms of the package. If a package has many thousands of lines—such as BoTorch—it is intuitive that it is more complex and thus more challenging to fully comprehend than a package with only a few hundred lines of code—such as NUBO. The number of code lines it takes to apply Bayesian optimization is less informative as it can easily be biased and distorted. Consider, for example, a very complex algorithm with many lines of code. It would be possible to wrap this algorithm into one function that can be called with one line of code. While this reduces the lines of code, it does not change the algorithm’s complexity. Thus, the comparison in this article focuses on the number of lines of the underlying code bases to give an idea of the size and complexity of the packages.
@@
Table 1: Overview of available Bayesian optimization packages in Python. We compare whether individual packages have a modular design and support sequential single-point, parallel multi-point, and asynchronous optimization. We also list the number of lines of code of the core package (without comments, examples, tests, etc.) and the version number.
@@
Although it is difficult to provide an exhaustive comparison of the relative efficiency of each of the packages, we have undertaken a limited comparison [^3] All comparisons were run on an Apple Mac mini with an M2 chip and 16 GB memory. of the following form. We compare NUBO to four of the packages mentioned above—BoTorch, bayes_opt, SMAC3 and pyGPGO—representing a reasonably wide range of complexity. All methods use Gaussian processes (introduced in Section 2.1) as the surrogate model and upper confidence bound (introduced in Section 2.2) as the acquisition function. Please see the replication materials published alongside this article for further details on the algorithms and the benchmarking. Two synthetic test functions from [24] were selected to benchmark the five packages. The first row of plots in Figure 1 compares the performance of sequential single-point optimization on A) the two-dimensional Levy function and B) the six-dimensional Hartmann function. The second row of plots compares the performance of parallel multi-point optimization with batches of four points on C) the two-dimensional Levy function and D) the six-dimensional Hartmann function. All functions are negated to transform them from their initial minimization problem into a maximization problem in line with the convention of Bayesian optimization. The plots provide the best observation at the current evaluation, averaging over ten replication runs. The results show that all packages converge towards the global optimum of 0.00 for the Levy and 3.32 for the Hartmann function. While NUBO requires more evaluations to find the global optimum for the Hartmann function (B) and D)) than the more complex packages such as BoTorch, it gets closest to the true optimum in all cases after all evaluations and shows low variance in these results (Table 2). These results show that the simplicity of NUBO’s implementation does not come at a cost in performance. NUBO can outperform packages with a similar level of complexity, such as pyGPGO and bayes_opt, and compares well against more complex packages, such as BoTorch and SMAC3. This is not to say that NUBO is the superior package for any problem, but rather that NUBO performs competitively while focusing on a transparent and simple design. This makes NUBO a good candidate for optimizing expensive black-box functions in the sciences—such as physical experiments and computer simulators—where transparency is vital.
@@
Figure 1: Comparison of different Python packages for Bayesian optimization. A) Sequential single-point optimization on the 2D Levy function; B) Sequential single-point optimization on the 6D Hartmann function; C) Parallel multi-point optimization with a batch size of four on the 2D Levy function; D) Parallel multi-point optimization with a batch size of four on the 6D Hartmann function.
@@
Table 2: Comparison of different Python packages for Bayesian optimization. The best observations averaged across the ten runs with corresponding standard errors are given for each package.
However, the time NUBO requires to complete one iteration with a maximum of 2.20s for D) is, on average, higher than for the other packages (Table 3). While this might be important for some areas of optimization, it is negligible when optimizing expensive black-box functions, as these functions are much more resource-intensive to evaluate. Thus, the small number of additional seconds that NUBO requires per iteration is insignificant compared to the resources required to conduct an experiment or a simulation.
@@
Table 3: Comparison of different Python packages for Bayesian optimization. The elapsed time per iteration averaged across the ten runs is given for each package.
Besides implementations in Python, there are also some implementations in other programming languages. For example, rBayesianOptimization [25] and ParBayesianOptimization [26] implement basic Bayesian optimization algorithms for hyper-parameter tuning similar to bayes_opt and pyGPGO in R. ParBayesianOptimization provides additional support for parallel optimization and follows [27].
The remainder of this paper is structured as follows. In Section 2, we introduce the Bayesian optimization algorithm, including Gaussian processes that form the surrogate model and acquisition functions that guide the optimization. The implementation of Bayesian optimization in NUBO is discussed in Section 3 before we illustrate how NUBO can be used to optimize expensive black-box functions through a non-trivial case study in Section 4. Finally, we draw conclusions and give an outlook on future work in Section 5.
2 Bayesian optimization
Bayesian optimization aims to solve the $d$-dimensional maximization problem
@@ 2 Bayesian optimization
where the input space is usually continuous and bounded by a hyper-rectangle $\mathcal{X}\in[a,b]^{d}$ with $a,b\in\mathbb{R}$. The function $f(\boldsymbol{x})$ is most commonly a derivative-free, expensive-to-evaluate black-box function that allows inputs $\boldsymbol{x}_{i}$ to be queried and outputs $y_{i}$ to be observed without gaining any further insights into the underlying system [1]. We assume any noise $\epsilon$ introduced when taking measurements to be independent and identically distributed Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ such that $y_{i}=f(\boldsymbol{x}_{i})+\epsilon$. Hence, a set of $n$ pairs of input data points and corresponding observations is defined as
$\mathcal{D}_{n}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}.$ (2)
@@
We further define training inputs as the matrix $\boldsymbol{X}_{n}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ and their training outputs as the vector $\boldsymbol{y}_{n}=\{y_{i}\}_{i=1}^{n}$. Simulators and experiments in various disciplines can be formulated to fit this description, including but not limited to the examples given in the introduction.
@@
Bayesian optimization [1, 28, 9, 11, 10] is a surrogate model-based optimization algorithm that aims to maximize the objective function $f(\boldsymbol{x})$ in a minimum number of function evaluations. Typically, the objective function does not have a known or analytical solvable mathematical expression, and every function evaluation is expensive. Such problems require a cost-effective and sample-efficient optimization strategy. Bayesian optimization meets these criteria by representing the objective function through a surrogate model $\mathcal{M}$, often a Gaussian process. This representation can be used to find the input points to be evaluated sequentially by maximizing a criterion specified through an acquisition function $\alpha(\cdot)$. A popular criterion is the upper confidence bound (UCB). This acquisition function can be classed as an optimistic acquisition function that considers the upper bound of the uncertainty around the surrogate model’s prediction to be true [11]. Bayesian optimization is performed in a loop, where training data is used to fit the surrogate model before the next point suggested by the acquisition function is evaluated and added to the training data (see the Algorithm 1 below). The process then restarts and gathers more information about the objective function with each iteration. Bayesian optimization is run for as many iterations as the evaluation budget $N$ allows, as shown in Algorithm 1, until a satisfactory solution is found or until a predefined stopping criterion is met.
Algorithm 1 Bayesian optimization algorithm
@@
Evaluation budget $N$, number of initial points $n_{0}$, surrogate model $\mathcal{M}$, acquisition function $\alpha$.
@@
Sample $n_{0}$ initial training data points $\boldsymbol{X}_{0}$ via a space-filling design [29] and gather observations $\boldsymbol{y}_{0}$.
@@
     Find $\boldsymbol{x}_{n}^{*}$ that maximizes an acquisition criterion $\alpha$ based on model $\mathcal{M}$.
@@
Figure 2 illustrates how the Bayesian optimization algorithm works for an optimization loop that runs for eight iterations and starts with two initial training points. In this example, NUBO finds an approximation of the global optimum ($x=8$) for a simple 1-dimensional problem on iteration seven. The surrogate model uses the available observations to provide a prediction and associated uncertainty (here shown as 95% confidence intervals around the prediction). This is our best estimate of the underlying objective function. This estimate is then used in the acquisition function to evaluate which input value will likely return a high output. Maximizing the acquisition function provides the next candidate point to be observed from the objective function before it is added to the training data and the whole process is repeated. Figure 2 shows how the surrogate model converges to the true objective function with each iteration. The acquisition function covers the input space by exploring regions with high uncertainty and exploiting regions with a high prediction. This property, the exploration-exploitation trade-off, is a cornerstone of the acquisition functions provided in NUBO.
Figure 2: Bayesian optimization applied to a 1-dimensional function with one local and one global maximum. Upper confidence bound is used as the acquisition function. The input space is bounded by $[0,10]$.

### v3

# Mike Diessner , Kevin J. Wilson

@@ Abstract
NUBO, short for Newcastle University Bayesian Optimisation, is a Bayesian optimisation framework for the optimisation of expensive-to-evaluate black-box functions, such as physical experiments and computer simulators. Bayesian optimisation is a cost-efficient optimisation strategy that uses surrogate modelling via Gaussian processes to represent an objective function and acquisition functions to guide the selection of candidate points to approximate the global optimum of the objective function. NUBO itself focuses on transparency and user experience to make Bayesian optimisation easily accessible to researchers from all disciplines. Clean and understandable code, precise references, and thorough documentation ensure transparency, while user experience is ensured by a modular and flexible design, easy-to-write syntax, and careful selection of Bayesian optimisation algorithms. NUBO allows users to tailor Bayesian optimisation to their specific problem by writing the optimisation loop themselves using the provided building blocks. It supports sequential single-point, parallel multi-point, and asynchronous optimisation of bounded, constrained, and/or mixed (discrete and continuous) parameter input spaces. Only algorithms and methods that are extensively tested and validated to perform well are included in NUBO. This ensures that the package remains compact and does not overwhelm the user with an unnecessarily large number of options. The package is written in Python but does not require expert knowledge of Python to optimise your simulators and experiments. NUBO is distributed as open-source software under the BSD 3-Clause licence.
Keywords Bayesian optimisation $\cdot$ black-box optimisation $\cdot$ surrogate model $\cdot$ Gaussian process $\cdot$ Monte Carlo $\cdot$ design of experiments $\cdot$ Python
@@ 1 Introduction
The optimisation of expensive black-box functions is a common problem encountered by researchers in a wide range of disciplines, such as engineering, computing, and natural sciences. These functions are characterised by an unknown or not analytically solvable mathematical expression and high costs of evaluation. The principal way to gather information about a black-box function is to provide it with some inputs and observe its corresponding output. However, this process produces high costs, for example, material costs for physical experiments, computing costs for simulators, or time costs in general (Frazier, 2018). Many optimisation algorithms, such as Adam (Kingma and Ba, 2014), L-BFGS-B (Zhu et al., 1997), and differential evolution (Storn and Price, 1997), rely either on derivative information about the objective function or large numbers of function evaluations. Neither is typically feasible when working with an expensive black-box function, requiring us to search elsewhere for a cost-effective and sample-efficient alternative.
@@
Bayesian optimisation takes a surrogate model-based approach with the aim of optimising expensive black-box functions in a minimum number of function evaluations. Although the genesis of Bayesian optimisation can be traced back to the middle of the 20th century (Kushner, 1964; Žilinskas, 1975; Močkus, 1975, 1989), it gained considerable popularity in the last two decades (Jones et al., 1998; Snoek et al., 2012; Shahriari et al., 2015; Frazier, 2018). In recent years, it has been applied to simulators and experiments in various research areas. For example, Bayesian optimisation was used in the field of computational fluid dynamics to maximise the drag reduction via the active control of blowing actuators (Diessner et al., 2022; O’Connor et al., 2023; Mahfoze et al., 2019), in chemical engineering for molecular design, drug discovery, molecular modelling, electrolyte design, and additive manufacturing (Wang and Dowling, 2022), and in computer science to fine-tune hyper-parameters of machine learning models (Wu et al., 2019) and for architecture search of neural networks (White et al., 2021).
@@
With NUBO, we provide an open-source implementation of Bayesian optimisation aimed at researchers with expertise in disciplines other than statistics and computer science. To ensure that our target audience can understand and use Bayesian optimisation to its full potential, NUBO focuses particularly on (a) transparency through clean and understandable code, precise references, and thorough documentation and (b) user experience through a modular and flexible design, easy syntax, and a careful selection of implemented algorithms. Various Python packages for Bayesian optimisation exist as listed in Table 1. Most of them only support sequential single-point optimisation, i.e., every point suggested by the algorithm has to be evaluated by the objective function before moving on to the next iteration. However, in many cases, parallelism can be exploited to speed up the optimisation process. For example, consider a simulator that can be run in parallel. Evaluating all points in parallel would save time as it would only take as long as evaluating a single point sequentially. pyGPGO (Jiménez and Ginebra, 2017), bayes_opt [^1] The package is also known under the name bayesian-optimization. (Nogueira, 2014), Spearmint (Havard University et al., 2014), and SMAC3 (Lindauer et al., 2022) do not allow parallel multi-point optimisation. Furthermore, Spearmint is not modular, resulting in less flexible implementations and giving the user less control when tailoring Bayesian optimisation to unique research problems. To our knowledge, the closest available package to NUBO is BoTorch (Balandat et al., 2020) as it also supports parallel and asynchronous optimisation through the use of Monte Carlo approximations of the acquisition functions. However, compared to the lightweight implementation of NUBO, BoTorch uses a very large code base that makes code comprehension difficult, as it often requires retracing various functions and objects through a large number of files. This can be quantified by the huge codebase represented in Table 1 as the total number of lines of code [^2] The total number of lines of code does not include comments, blank lines and files that are irrelevant to the actual algorithms, such as examples, tests and test functions.: NUBO implements Bayesian optimisation in only 1,322 lines of code over 20 files, while BoTorch uses 38,419 lines of code—roughly 29 times more than NUBO—and spreads them between 160 files. It also provides a large number of functions and methods that enforce decisions non-expert users do not have the knowledge and experience to make. NUBO lightens this burden of the user by limiting itself to the most important methods. Table 1 also includes GPyOpt (The GPyOpt authors, 2016); however, it is no longer maintained and has recently been archived.
@@
The number of code lines regards the underlying code bases of the packages and not the lines of code a user must write to apply Bayesian optimisation. When talking about a transparent implementation, the former is a better proxy as it reflects the complexity of the whole package, that is, all functions and algorithms of the package. If a package has many thousands of lines—such as BoTorch—it is intuitive that it is more complex and thus harder to fully comprehend than a package with only a few hundred lines of code—such as NUBO. The number of code lines it takes to apply Bayesian optimisation is less informative as it can easily be biased and distorted. Consider, for example, a very complex algorithm with many lines of code. It would be possible to wrap this algorithm into one function that can be called with one line of code. While this reduces the lines of code, it does not change the algorithm’s complexity. Thus, the comparison in this article focuses on the number of lines of the underlying code bases to give an idea of the size and complexity of the packages.
@@
Table 1: Overview of available Bayesian optimisation packages in Python. We compare whether individual packages have a modular design and support sequential single-point, parallel multi-point, and asynchronous optimisation. We also list the number of lines of code of the core package (without comments, examples, tests, etc.) and the version number.
@@
Although it is difficult to provide an exhaustive comparison of the relative efficiency of each of the packages, we have undertaken a limited comparison [^3] All comparisons were run on an Apple Mac mini with a M2 chip and 16 GB memory. of the following form. We compare NUBO to four of the packages mentioned above—BoTorch, bayes_opt, SMAC3 and pyGPGO—representing a reasonably wide range of complexity. All methods use Gaussian processes (introduced in Section 2.1) as the surrogate model and upper confidence bound (introduced in Section 2.2) as the acquisition function. Please see the replication materials published alongside this article for further details on the algorithms and the benchmarking. Two synthetic test functions from Surjanovic and Bingham were selected to benchmark the five packages. The first row of plots in Figure 1 compares the performance of sequential single-point optimisation on A) the two-dimensional Levy function and B) the six-dimensional Hartmann function. The second row of plots compares the performance of parallel multi-point optimisation with batches of four points on C) the two-dimensional Levy function and D) the six-dimensional Hartmann function. All functions are negated to transform them from their initial minimisation problem into a maximisation problem in line with the convention of Bayesian optimisation. The plots provide the best observation at the current evaluation, averaging over ten replication runs. The results show that all packages converge towards the global optimum of 0.00 for the Levy and 3.32 for the Hartmann function. While NUBO requires more evaluations to find the global optimum for the Hartmann function (B) and D)) than the more complex packages such as BoTorch, it gets closest to the true optimum in all cases after all evaluations and shows low variance in these results (Table 2). These results show that the simplicity of NUBO’s implementation does not come at a cost in performance. NUBO can outperform packages with a similar level of complexity, such as pyGPGO and bayes_opt, and compares well against more complex packages, such as BoTorch and SMAC3. This is not to say that NUBO is the superior package for any problem, but rather that NUBO performs competitively while focusing on a transparent and simple design. This makes NUBO a good candidate for the optimisation of expensive black-box functions in the sciences—such as physical experiments and computer simulators—where transparency is vital.
@@
Figure 1: Comparison of different Python packages for Bayesian optimisation. A) Sequential single-point optimisation on the 2D Levy function; B) Sequential single-point optimisation on the 6D Hartmann function; C) Parallel multi-point optimisation with a batch size of four on the 2D Levy function; D) Parallel multi-point optimisation with a batch size of four on the 6D Hartmann function.
@@
Table 2: Comparison of different Python packages for Bayesian optimisation. The best observations averaged across the ten runs with corresponding standard errors are given for each package.
However, the time NUBO requires to complete one iteration with a maximum of 2.20s for D) is, on average, higher than for the other packages (Table 3). While this might be important for some areas of optimisation, it is negligible when it comes to the optimisation of expensive black-box functions, as these functions are much more resource-intensive to evaluate. Thus, the small number of additional seconds that NUBO requires per iteration is insignificant compared to the resources required to conduct an experiment or a simulation.
@@
Table 3: Comparison of different Python packages for Bayesian optimisation. The elapsed time per iteration averaged across the ten runs is given for each package.
Besides implementations in Python, there are also some implementations in other programming languages. For example, rBayesianOptimization (Yan, 2024) and ParBayesianOptimization (Wilson, 2022) implement basic Bayesian optimisation algorithms for hyper-parameter tuning similar to bayes_opt and pyGPGO in R. ParBayesianOptimization provides additional support for parallel optimisation and follows Wilson et al. (2018).
The remainder of this paper is structured as follows. In Section 2, we introduce the Bayesian optimisation algorithm, including Gaussian processes that form the surrogate model and acquisition functions that guide the optimisation. The implementation of Bayesian optimisation in NUBO is discussed in Section 3 before we illustrate how NUBO can be used to optimise expensive black-box functions through a non-trivial case study in Section 4. Finally, we draw conclusions and give an outlook on future work in Section 5.
2 Bayesian optimisation
Bayesian optimisation aims to solve the $d$-dimensional maximisation problem
@@ 2 Bayesian optimisation
where the input space is usually continuous and bounded by a hyper-rectangle $\mathcal{X}\in[a,b]^{d}$ with $a,b\in\mathbb{R}$. The function $f(\boldsymbol{x})$ is most commonly a derivative-free, expensive-to-evaluate black-box function that allows inputs $\boldsymbol{x}_{i}$ to be queried and outputs $y_{i}$ to be observed without gaining any further insights into the underlying system (Frazier, 2018). We assume any noise $\epsilon$ introduced when taking measurements to be independent and identically distributed Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2})$ such that $y_{i}=f(\boldsymbol{x}_{i})+\epsilon$. Hence, a set of $n$ pairs of input data points and corresponding observations is defined as
$\mathcal{D}_{n}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$
@@
and we further define training inputs as the matrix $\boldsymbol{X}_{n}=\{\boldsymbol{x}_{i}\}_{i=1}^{n}$ and their training outputs as the vector $\boldsymbol{y}_{n}=\{y_{i}\}_{i=1}^{n}$. Simulators and experiments in various disciplines can be formulated to fit this description, including but not limited to the examples given in the introduction.
@@
Bayesian optimisation (Frazier, 2018; Gramacy, 2020; Jones et al., 1998; Shahriari et al., 2015; Snoek et al., 2012) is a surrogate model-based optimisation algorithm that aims to maximise the objective function $f(\boldsymbol{x})$ in a minimum number of function evaluations. Typically, the objective function does not have a known or analytical solvable mathematical expression, and every function evaluation is expensive. Such problems require a cost-effective and sample-efficient optimisation strategy. Bayesian optimisation meets these criteria by representing the objective function through a surrogate model $\mathcal{M}$, often a Gaussian process. This representation can be used to find the input points to be evaluated sequentially by maximising a criterion specified through an acquisition function $\alpha(\cdot)$. A popular criterion is the upper confidence bound (UCB). This acquisition function can be classed as an optimistic acquisition function that considers the upper bound of the uncertainty around the surrogate model’s prediction to be true (Shahriari et al., 2015). Bayesian optimisation is performed in a loop, where training data is used to fit the surrogate model before the next point suggested by the acquisition function is evaluated and added to the training data (see the Algorithm 1 below). The process then restarts and gathers more information about the objective function with each iteration. Bayesian optimisation is run for as many iterations as the evaluation budget $N$ allows, as shown in Algorithm 1, until a satisfactory solution is found or until a predefined stopping criterion is met.
Algorithm 1 Bayesian optimisation algorithm
@@
Evaluation budget NN, number of initial points $n_{0}$, surrogate model $\mathcal{M}$, acquisition function $\alpha$.
@@
Sample $n_{0}$ initial training data points $\boldsymbol{X}_{0}$ via a space- filling design (McKay et al., 1979) and gather observations $\boldsymbol{y}_{0}$.
@@
  Find $\boldsymbol{x}_{n}^{*}$ that maximises an acquisition criterion $\alpha$ based on model $\mathcal{M}$.
@@
Figure 2 illustrates how the Bayesian optimisation algorithm works for an optimisation loop that runs for eight iterations and starts with two initial training points. In this example, NUBO finds an approximation of the global optimum ($x=8$) for a simple 1-dimensional problem on iteration seven. The surrogate model uses the available observations to provide a prediction and associated uncertainty (here shown as 95% confidence intervals around the prediction). This is our best estimate of the underlying objective function. This estimate is then used in the acquisition function to evaluate which input value is likely to return a high output. Maximising the acquisition function provides the next candidate point to be observed from the objective function before it is added to the training data and the whole process is repeated. Figure 2 shows how the surrogate model converges to the true objective function with each iteration. The acquisition function covers the input space by exploring regions with high uncertainty and exploiting regions with a high prediction. This property, known as the exploration-exploitation trade-off, is a cornerstone of the acquisition functions provided in NUBO.
Figure 2: Bayesian optimisation applied to a 1-dimensional function with one local and one global maximum. Upper confidence bound is used as the acquisition function. The input space is bounded by $[0,10]$.